OpenStudy (anonymous):

Is this right? i was trying to simplify... (4-c)/(2c-8)=(4-c)/(c∙c-8)=4/(c-8)=1 ⁄ (c-4)

OpenStudy (anonymous):

@amistre64?

OpenStudy (amistre64):

2c does not equal c*c 2c = c+c but the key here is to factor out a -2 from the bottom, then cancel

OpenStudy (amistre64):

\[\frac{(4-c)}{(2c-8)}\] \[\frac{(4-c)}{-2(-c+4)}\] \[\frac{\cancel{(4-c)}}{-2\cancel{(4-c)}}\]

OpenStudy (anonymous):

so 1/-2?